# AI-boosted rare event sampling to characterize extreme weather

### v1

@@ Abstract
Assessing the frequency and intensity of extreme weather events, and understanding how climate change affects them, is crucial for developing effective adaptation and mitigation strategies [1, 2, 3, 4, 5]. However, observational datasets are too short and physics-based global climate models (GCMs) are too computationally expensive to obtain robust statistics for the rarest, yet most impactful, extreme events [6]. AI-based emulators have shown promise for predictions at weather and even climate timescales [7, 8, 9, 10, 11, 12, 13], but they struggle on extreme events with few, or no, examples in their training dataset [14, 15, 16, 17]. Rare event sampling (RES) algorithms have previously demonstrated success for some extreme events [18, 19, 20, 21, 22, 23], but their performance depends critically on a hard-to-identify “score function”, which guides efficient sampling by a GCM. Here, we develop a novel algorithm, AI+RES, which uses ensemble forecasts of an AI weather emulator [7, 8, 9, 10] as the score function to guide highly efficient resampling of the GCM and generate robust (physics-based) extreme weather statistics and associated dynamics at $30-300\times$ lower cost. We demonstrate AI+RES on mid-latitude heatwaves, a challenging test case requiring a score function with predictive skill many days in advance. AI+RES, which synergistically integrates AI, RES, and GCMs, offers a powerful, scalable tool for studying extreme events in climate science, as well as other disciplines in science and engineering where rare events and AI emulators are active areas of research.
Introduction
Rare, extreme weather and climate events have enormous economic and social costs [24, 25], but it is difficult to estimate their frequency and potential impact from observational data or high-fidelity physics-based GCM simulations. Extreme value theory [26] can be used to extrapolate from short historical records, but it is often overwhelmed by uncertainty [27, 28, 6, 22], and does not provide examples of the event of interest for physical interrogation. GCMs require simulation lengths at least ten times longer than the return time of the event of interest for accurate sampling, which is not feasible given the computational cost of state-of-the-art physics-based GCMs to fully resolve all relevant scales and processes. In response to this need, rare event sampling (RES) and artificial intelligence (AI) emulator techniques have been developed, but both have proven insufficient on their own to fully resolve the problem.
RES is a set of importance sampling tools that focus computation on the rare event of interest [18, 19, 29, 20, 30, 22, 23]. RES entails scheduled duplication (“splitting”) or termination (“killing”) of members of an ensemble of model simulations, that otherwise evolve independently, to promote efficient sampling of the targeted rare event (Fig. 1a). The probability that any one ensemble member (a trajectory) is duplicated or terminated is determined by a user-chosen score function that can be evaluated for any trajectory at the scheduled resampling times. Despite promising results, the full power of RES has not been realized for climate and weather due primarily to the difficulty in identifying effective score functions. Choosing a score function typically requires extensive domain knowledge and a process of trial and error with associated costs in computation and time that may swamp any advantages of using RES. Standard RES uses simple persistence as a score function: it relies on the current value of the index variable (which defines the extreme event of interest; e.g., 5-day mean temperature), rather than its value at the time of the event. This choice of score function has proven highly effective for sampling extremes of long-term averages, e.g., seasonal means [18, 19, 21, 22, 23].
However, this approach fails catastrophically when the current value of the index variable is a poor predictor of its future value [19, 20], as is the case for many phenomena in complex atmospheric dynamics; examples include blocking-driven heatwaves, which have led to some of the most socio-economically impactful extreme events in recent memory in France, Russia, and U.S. Pacific Northwest [31, 32, 33]. Previous work has therefore established the strong potential of RES, but only if an efficient method for finding an effective score function can be identified. Recently, a set of interesting papers studied unprecedented extreme events using ensemble boosting, a cousin of RES [34, 35]. In its initial formulation [34], ensemble boosting could not be used to estimate event probabilities. In a more recent version [35, 36], however, probabilities were estimated by adopting a rare event simulation approach with a one-step splitting scheme. This similar method could benefit from an analogous AI score function.
One of the most exciting recent developments in climate science and scientific AI has been the introduction of auto-regressive AI weather emulators such as FourCastNet, Pangu, GraphCast, and GenCast [7, 8, 9, 10]. AI weather emulators are neural networks that are trained on spatio-temporal data from historical observations or physics-based GCM simulations to advance the state of the system forward in time for a short period (e.g., 6 hours) and then applied recursively to produce forecasts over longer periods. Various analyses have confirmed that AI weather emulators trained on high-resolution observation-derived reanalysis datasets can outperform the best physics-based numerical weather models for short- and medium-range ($\sim$10-15 day) forecasts [9, 8, 37]. More importantly, these emulators are up to $10^{4}$ times faster than state-of-the-art physics-based models [7], suggesting that they might be useful in reducing the uncertainty associated with rare extreme weather events. Specifically, extremely long and/or large ensemble emulations can be generated so that accurate “direct sampling” is possible, even for the rarest events [38, 39]. The problem with this method, however, is that AI emulators are trained on available historical records or expensive high-fidelity GCM simulations that may contain few, or no, examples of the rarest extreme events. Characterizing rare events using direct simulation with an AI emulator would therefore require not only reliable out-of-distribution extrapolation to extreme events beyond the training dataset but also capturing their frequency correctly, which recent work has suggested the AI emulators struggle with [14, 15, 17].
In this paper, we leverage advances in AI weather emulation to build an effective score function for RES, combining recent developments in both areas to construct a robust algorithm to characterize rare, extreme events (Fig. 1). As a demonstration of this new framework (AI+RES), we focus on simulating extreme heat waves and apply our approach to a GCM of intermediate complexity, PlaSim [40], which allows for extensive long-term control runs to compute baseline rare event statistics and rigorously evaluate the performance of our approach. Heat waves are a well-motivated application area because they are the deadliest extreme weather events and are expected to worsen under climate change [41, 42, 43]. We show that AI+RES provides accurate rare event statistics at a numerical speed-up of up to a factor of $\mathcal{O}(100)$, whereas standard RES fails to even yield examples of the rarest events of interest. This paper serves as a demonstration of a novel methodology that can overcome the primary difficulty of RES and unleash its power widely across any field for which an AI emulator can be constructed.
@@ Introduction
Figure 1: Schematic of the AI+RES framework. We train an AI weather emulator on 100 years of PlaSim GCM data, then use it to guide a RES algorithm. (a) In our proposed method (AI+RES), we run a PlaSim ensemble simulation with NN parallel simulations, called walkers and denoted $X_{t}^{i}$. (b) At each resampling time $t_{k}$ in the algorithm (vertical dashed lines), we perform an ensemble forecast with the AI emulator for each PlaSim walker until the target time $t_{f}$, then duplicate the more promising ones based on these forecasts (Algorithm 1 and Eq. (6)). This allows us to both generate a large catalog of extreme events from physics-based simulations (c) and to compute unbiased probabilities for these rare events (Eq. (4)) (d) at a fraction of the cost of Direct Numerical Simulation (DNS). An example of schematic (a) with actual data from an AI+RES experiment can be found in Extended Data Fig. 3.
AI+RES
Figure 1 outlines the AI+RES algorithm. We use an AI emulator, trained to predict the state of a physical system at time $t+\Delta t$ conditioned on the state at time $t$, to address the major shortcoming in RES algorithms, namely, the choice of a score function $\theta$.
The RES splitting algorithm proceeds as follows: we perform NN parallel simulations (walkers) using a physics-based model (e.g., a GCM), starting at time $t_{0}$ with small initial condition perturbations. At user-specified resampling times, $\{t_{k}\}_{k=1}^{K}$, we selectively duplicate or terminate the $i^{th}$ walker according to the value of the score function $\theta^{i}_{k}$ until the walkers reach the final time horizon $t_{f}=t_{K}$. Walkers with a large score are more likely to be duplicated, while walkers with a small score are more likely to be terminated. After resampling, we continue the GCM simulation of the selected walkers until the next resampling time, when the process is repeated. Each walker is accompanied by a statistical weight $w^{i}_{k}$ that allows unbiased estimation of rare event statistics of interest.
RES applications have most often used simple persistence as the score function: the current value of the index variable (which defines the extreme event of interest, e.g., temperature) is used as a proxy for its future value. In this paper, we introduce a new approach for constructing the score function based on an AI weather emulator. At each resampling time $t_{k}$, and from every current walker state, we perform an ensemble forecast of the system state (with $M$ members) until the final time $t_{f}$ using the AI emulator (see the AI Emulator section in Methods). The score function for each RES walker is then the mean value of the index variable at time $t_{f}$ over the AI emulator ensemble forecast for that walker. To the extent that the AI emulator faithfully reproduces the GCM’s dynamics, the prediction furnished by the AI emulator ensemble is ideally suited for use in constructing the score function. Our results show that the AI emulator’s weather forecast skill is sufficient to provide a very effective score function for AI+RES. More details on our implementation are given in the Methods section, including our choice of hyperparameters for the RES Algorithm 1 in Extended Data Table 2.
@@
In this paper, we focus on the sampling of rare mid-latitude heatwaves due to their societal impact and the challenge they pose to traditional (standard) RES. Specifically, following previous studies [18, 44], we focus on events characterized by large values of the spatio-temporal average of 2-meter temperature T2m over a geographic region $\mathcal{R}$ during a time interval $[t,t+L]$. This evolving spatio-temporal average is our index variable $A_{L}(t)$:
$A_{L}(t):=\frac{1}{L}\int_{t}^{t+L}\left(\frac{1}{\mathcal{R}}\int_{\mathcal{R}}\text{T}_{2m}(\vec{r},s),d\vec{r}\right)ds.$ (1)
We want to estimate the return times of the events $A_{L}(t_{f})>a$ for large thresholds $a$. We use $L=7$ days, a window size small enough to capture the peak of mid-latitude heatwaves yet large enough to have significant societal impacts. In this study, we consider two regions, $\mathcal{R}$, each defined as a $3\times 3$ grid point domain centered over France and Chicago (Extended Data Fig. 1). We initialize simulations at $t_{0}$=July 2 and integrate forward the PlaSim GCM until $t_{f}+L$, with $t_{f}=$ August 1, chosen to coincide with the climatological peak of $A_{L}(t_{f})$.
@@
To evaluate the performance of AI+RES, we compare it to a ground truth and five baselines. The ground truth, direct numerical simulation (DNS) with $N=50,000$, is a very large ensemble that is only possible due to the computational efficiency of PlaSim. The first baseline, DNS, $N=400$, is a DNS with PlaSim using the same ensemble size as in the AI+RES algorithm. The second baseline, Standard-RES, uses the RES algorithm but with the traditional persistence score function used in state-of-the-art weather and climate applications [18, 19, 21, 22, 23]. The third baseline, AI-DNS, is a DNS ensemble using the AI emulator started from the same initial conditions as DNS and RES, with the same ensemble size as the ground truth ($N=50,000$). This allows us to determine whether the AI emulator has learned enough from its 100-year training period to extrapolate to tail events in PlaSim. The fourth baseline, EVT, applies Extreme Value Theory, specifically the Generalized Pareto Distribution (GPD) within a Peak-over-Threshold (POT) framework, to PlaSim datasets of the same size as the AI+RES experiments ($N=400$). We estimate uncertainty in this method by performing the fits on 100 different datasets of equal size. The fifth baseline, PFS+RES, yields an upper bound on algorithm performance using a perfect-forecast system as the score function. In this baseline, we use an ensemble forecast generated with PlaSim itself as the score function (Extended Data). However, this baseline is too expensive for more complex GCMs; it is included to show how the algorithm would perform with a “perfect emulator”.
@@
Figure 2: Return-period curve for the T2m 7-day average over France (left panel) and Chicago (right panel). Black dots are obtained with a $N=50,000$ member DNS and serve as ground truth. The red dots are from DNS, but with the same computational budget as AI+RES ($N=400$). The cyan dots are obtained by running a $N=50,000$ member ensemble with the AI emulator only (AI-DNS). The solid blue line shows the median return-period curve produced by 10 independent realizations of the AI+RES algorithm with $N=400$ walkers and the blue shaded area shows the 10th and 90th percentiles. The solid yellow line shows the median return-period curve produced by 10 independent realizations of the standard RES algorithm with $N=400$ walkers, and the yellow shaded area shows the 10th and 90th percentiles. The gray shaded area (EVT) is obtained by fitting different GPD distributions with 100 independent training datasets of size $N=400$, and showing the 10th and 90th percentiles as confidence interval.
In both France and Chicago, AI+RES produces accurate, unbiased return period estimates up to 50,000 years, despite using ensembles with only $N=400$ walkers; see Fig. 2 (the 50,000-year limit is imposed by the size of our reference ensemble for validating the procedure). In stark contrast, DNS of the same computational cost ($N=400$) can only produce accurate return period estimates up to about 50 years, 1,000 times shorter. This dramatic difference underscores the utility of RES when an effective score function can be identified. Moreover, Standard-RES saturates for return periods longer than about 100 years, falsely suggesting an upper limit on heatwave intensity. This indicates that persistence misses important trajectories to rare heat waves that the AI-based score function is able to identify effectively. The AI-DNS baseline is biased both in the mean and variability of the observable, producing return period estimates that fail even on a decadal timescale (the correct AI-DNS estimate at a return period of about 2,000 years in France is an accident of the curves crossing). It is notable that despite this failure of AI-DNS, the AI emulator’s weather forecast skill can still serve as a useful score function by effectively ranking progress of walkers toward strong heat waves. Finally, the correct return periods do fall within the uncertainty of the EVT baseline even at the longest timescales, but the uncertainty estimate is up to four times larger than that from AI+RES for return periods on the order of 10,000 years. The EVT estimates depend so strongly on the particular years fed into them that they cannot be relied on to produce accurate return period estimates. Also, EVT, unlike AI+RES, does not provide information about the identified extremes and their dynamics.
@@
Figure 3: Speed-up factors for the AI+RES algorithm over France and Chicago. Left panel: Variance speed-up factor as a function of return period (Eq. (13)). This measures the reduction in the variance of return period estimates compared to DNS (see Methods). Right panel: Sampling speed-up factor as a function of return period (Eq. (10)), averaged over 10 independent algorithm realizations. This measures the gain in the number of extreme samples produced by the algorithm compared to DNS (see Methods).
Now that we have established that AI+RES is unbiased, we can quantify its computational speed-up relative to DNS using efficiency at sampling extreme events (Eq. (10)) and variance of rare probability estimators (Eq. (13)). Both methods tend to yield greater computational gains for rarer events (Fig. 3), with speed-ups of more than 100 for a return period of 10,000 years. The only exception is that the variance speed-up estimate for the France region peaks at a factor of 25 around the 2,500-year return period, then gradually decreases to 10 for the rarest events. This non-monotonic trend is most likely due to a sampling error (see Variance speed-up factor in the Methods section). Moreover, we hypothesize that the difference in variance speed-up for the two regions is due to regional differences in AI emulator weather forecast skill, potentially due to differences in soil moisture variability, as pointed out in Fig. S6 in the SI.
@@
Figure 4: Composite maps of heatwaves over France with return periods exceeding 100 years. Events are defined by Eq. (1) with $L=$ 3 days. The first row shows daily mean Z500 anomaly composites three days before the heatwave onset; the second row shows the $L$-day average Z500 anomaly composites during the heatwave; the third row shows the $L$-day average T2m anomaly composites during the heatwave. Left column: DNS with N=50, [^0] N=50{,}000 (ground truth). Middle column: results from the AI+RES algorithm with $N=400$. Right column: DNS with $N=400$. See Fig. S5 in the SIfor the same analysis but over Chicago.
In addition to greatly reduced variance in return period estimations, a major advantage of AI+RES over the EVT baseline is that it generates full physical trajectories leading to the rare event of interest. This is particularly valuable for investigating the dynamics of the rare event and its precursors [23].
In Fig. 4, we illustrate this capability by presenting composite maps associated with heatwaves over France with a return period longer than 100 years from the ground-truth DNS (N=50, [^0] N=50{,}000), samples generated with AI+RES ($N=400$), and samples generated from the DNS with the same ensemble ($N=400$). A strong precursor pattern emerges in the ground-truth DNS three days before the heatwave onset, marked by a pronounced positive 500-hPa geopotential height anomaly over the British Isles, flanked by a negative anomaly over the Labrador Sea. A distinct negative anomaly off the coast of Portugal suggests potential influence from a cut-off low, while a weaker anomaly is visible over Eastern Europe. At the heatwave peak, this synoptic pattern shifts: the high-pressure anomaly migrates eastward, the Labrador Sea anomaly propagates into the North Atlantic, and the Eastern European anomaly intensifies. The AI+RES composite shows similar patterns and captures the correct evolution, although it is slightly noisier than the ground truth due to fewer samples. In contrast, the composite from the DNS with the same computational budget as AI+RES is significantly noisier and contains incorrect physical patterns. For example, the 3-day-lagged Z500 anomaly north of Scandinavia in this composite is a spurious teleconnection likely due to sampling error.
Similarly, Fig. S5 in the SI shows that AI+RES generates physically realistic samples over Chicago that match those from ground truth, while DNS with the same computational budget produces samples with spurious features, obscuring the underlying dynamics.

### v2

@@ Abstract
Weather extremes pose major societal risks, especially in a changing climate, but due to their rarity, they are difficult to study using limited observations or complex climate models. We introduce AI+RES, a framework coupling fast AI weather forecasts with a high-fidelity physics model using a rare-event algorithm to efficiently characterize extremes. This approach enables the study of the statistics and physics of very rare events, such as once per millennium heatwaves at two orders-of-magnitude lower computational cost. AI+RES can be applied broadly across climate science and other fields concerned with rare events.
@@
Figure 1: Schematic of the AI+RES framework. We first train an AI weather emulator on 100 years of PlaSim GCM data, then use it to guide a Rare Event Sampling (RES) algorithm. (a) In AI+RES, we run a PlaSim ensemble simulation with NN parallel simulations, called walkers and denoted $X_{t}^{i}$. (b) At each resampling time $t_{k}$ in the algorithm (vertical dashed lines), we perform an ensemble forecast with the AI emulator for each PlaSim walker until the target time $t_{f}$, then duplicate the more promising walkers based on these forecasts (Algorithm 1 and Eq. (3)). This allows us to both generate a large catalog of extreme events from physics-based simulations (c) and to compute unbiased probabilities for these rare events (Eq. (2)) (d) at a fraction of the cost of Direct Numerical Simulation (DNS). An example of schematic (a) with actual data from an AI+RES experiment can be found in Supplemental Material (SM) Fig. S9.
Rare, extreme weather and climate events—such as heatwaves, floods, and hurricanes—have enormous socio-economic impacts [1, 2]. Accurately assessing how their frequency and intensity change in a warming climate is essential for developing effective mitigation and adaptation strategies [3, 4, 5, 6, 7]. However, it is difficult to estimate the frequency and potential impact of the rarest events from observational data [8]. Extreme value theory [9] can be used to extrapolate from short historical records, but it is often overwhelmed by uncertainty [10, 11, 8, 12], and does not provide examples of the event of interest for physical interrogation. High-fidelity physics-based global climate model (GCM) simulations can provide additional data, but accurate sampling requires simulation lengths at least ten times longer than the return time of the event of interest, which is computationally infeasible. In response to these computational constraints, rare event sampling (RES) techniques and artificial intelligence (AI) emulators have been developed. While promising, both methods have proven insufficient on their own to fully resolve the problem.
RES is a set of importance sampling tools that focus computation on the rare event of interest [13, 14, 15, 16, 17, 12, 18]. RES entails scheduled duplication (“splitting”) or termination (“killing”) of members of an ensemble of independently-evolving model simulations to promote efficient sampling of the targeted rare event. The probability that any one ensemble member (a trajectory) is duplicated or terminated is determined by a user-chosen score function evaluated at the scheduled resampling times. Choosing a score function typically requires extensive domain knowledge, trial and error, and parameter optimization with costs that may outweigh any advantages of using RES over direct simulation. Recent work has also explored using ensemble boosting, a cousin of RES, to study unprecedented extreme events [19, 20]. This approach has recently been extended to allow probability estimation by adopting an RES approach with a one-step splitting scheme [20, 21].
RES has proven highly effective for sampling extremes of long-duration weather, e.g., seasonal means [13, 14, 22, 12, 18], by using persistence, i.e., the current value of the index variable defining the event, as a score function. This approach, however, fails catastrophically when the current value of the index variable is a poor predictor of its future value [14, 16]. This is the case for many phenomena in complex atmospheric dynamics, such as blocking-driven heatwaves, which have led to some of the most socio-economically impactful extreme events in recent memory in France, Russia, and U.S. Pacific Northwest [23, 24, 25, 26]. Previous work has therefore established the strong potential of RES, but only if an efficient method for finding an effective and accessible score function can be identified.
One of the most exciting recent developments in climate science and scientific AI has been the introduction of auto-regressive AI weather emulators [27, 28, 29, 30]. When trained on high-resolution reanalysis data, these models can outperform state-of-the-art physics-based numerical weather prediction systems for short- and medium-range ($\sim$10–15 day) forecasts [29, 28, 31]. Their extreme computational efficiency (up to $10^{4}$ times faster than traditional models [27]) makes them appealing for studying rare extreme weather. More recently, stable AI emulators have been extended to climate timescales, enabling extremely long simulations or large ensembles in which rare events could, in principle, be sampled directly [32, 33, 34, 35, 36]. However, because these models are trained to predict at short lead times, and only on historical data or expensive high-fidelity simulations that contain few examples of the most extreme events, they often struggle to extrapolate reliably and to reproduce the correct frequencies of very rare extremes [37, 38, 39, 40].
In this paper, we present AI+RES, a robust algorithm that leverages advances in AI emulation to build an effective score function for RES to characterize rare, extreme events. This algorithm is described in Fig. 1 and discussed in detail in the End Matter. We demonstrate this new framework by using it to simulate extreme heat waves in an intermediate complexity GCM, PlaSim [41], which allows for extensive long-term control runs to compute baseline rare event statistics and rigorously evaluate the performance of our approach. Heat waves are a well-motivated application area because they are the deadliest extreme weather events and are expected to worsen under climate change [42, 43, 44]. We show that AI+RES provides accurate rare event statistics at a numerical speed-up of up to a factor of $\mathcal{O}(100)$, whereas standard RES fails to even yield examples of the rarest events of interest. This paper serves as a demonstration of a novel methodology that can overcome the primary difficulty of RES and unleash its power widely across any field for which a fast AI emulator can be constructed.
@@
As mentionned, we focus on sampling rare mid-latitude heatwaves due to their societal impact and the challenge they pose to standard RES. Following previous studies [13, 45], we consider heatwave events defined by large values of the spatio-temporal average of the 2-meter air temperature, T2m, over a prescribed geographic region and a time interval typically spanning one week during the summer season (see Eq.(1)). We consider two such regions, centered over France and over Chicago, respectively (SM Fig.S7).
@@
To evaluate the performance of AI+RES, we compare it to a ground truth and five baselines. The ground truth, direct numerical simulation (DNS) with $N=50,000$, is a very large ensemble that is only possible due to the computational efficiency of PlaSim. The first baseline, DNS, $N=400$, is a DNS with PlaSim using the same ensemble size as the number of walkers used in the AI+RES algorithm. The second baseline, Standard-RES, uses the RES algorithm but with the traditional persistence score function used in state-of-the-art weather and climate applications [13, 14, 22, 12, 18]. The third baseline, AI-DNS, is a DNS ensemble using the AI emulator started from the same initial conditions as DNS and RES, with the same ensemble size as the ground truth ($N=50,000$). This allows us to determine whether the AI emulator has learned enough from its 100-year training period to extrapolate to tail events in PlaSim. The fourth baseline, EVT, applies Extreme Value Theory, specifically the Generalized Pareto Distribution (GPD) within a Peak-over-Threshold (POT) framework, to PlaSim datasets of the same size as the AI+RES experiments ($N=400$). We estimate uncertainty in this method by performing the fits on 100 different datasets of equal size. We discuss a fifth baseline, PFS+RES, in SM Sec. S8.2. This baseline uses a PlaSim ensemble forecast for the score function and yields an upper bound on algorithm performance given a “perfect emulator”. This baseline is only possible with GCMs like PlaSim that are computationally inexpensive.
@@
Figure 2: Return-period curve for a 7-day heatwave (T2m averaged over 7-days) over France (left panel) and Chicago (right panel). The solid blue and yellow lines show the median return-period curve from 10 independent realizations of the AI+RES and standard-RES algorithm, respectively; shaded regions indicate the 10th–90th percentiles. Red and Black dots show DNS with N=50,000 and N=400. The gray shaded area (EVT) shows the 10th and 90th percentiles obtained by fitting different GPD distributions with 100 independent training datasets with N=400.
Figure 2 shows that for both France and Chicago, AI+RES produces accurate, unbiased return period estimates up to our reference ensemble limit of 50,000 years, despite using ensembles with only $N=400$ walkers. In stark contrast, DNS of the same computational cost ($N=400$) can only produce accurate return period estimates up to about 50 years, 1,000 times shorter. This dramatic difference underscores the utility of RES when an effective score function can be identified. Moreover, Standard-RES saturates for return periods longer than about 100 years, falsely suggesting an upper limit on heatwave intensity. This indicates that persistence misses important trajectories to rare heat waves that the AI-based score function is able to identify. The AI-DNS baseline is biased both in the mean and variability of the observable, producing erroneous return period estimates even on a decadal timescale. The correct AI-DNS estimate at a return period of about 2,000 years in France is an accident of the curves crossing. It is notable that despite the errors in AI-DNS, the AI emulator’s weather forecast skill can still serve as a useful score function by effectively ranking progress of walkers toward strong heat waves. Finally, the correct return periods do fall within the uncertainty of the EVT baseline even at the longest timescales, but the uncertainty estimate is up to four times larger than that from AI+RES for return periods on the order of 10,000 years. The EVT estimates depend so strongly on the particular years fed into them that they cannot be relied on to produce accurate return period estimates. Also, EVT, unlike AI+RES, does not provide information about the identified extremes and their dynamics.
@@
Figure 3: Speed-up factors for the AI+RES algorithm over France and Chicago as a function of return period for reducing variance in estimated probability and producing additional samples.
Having established that AI+RES is unbiased, we can quantify its computational speed-up compared to DNS using the relative efficiency of sampling extreme events (SM Eq. (S7)) and the relative variance of rare probability estimators (SM Eq. (S10)). Both metrics tend to yield greater computational gains for rarer events (Fig. 3), with speed-ups of more than 100 for a return period of 10,000 years. The only exception is that the variance speed-up estimate for the France region peaks at a factor of 25 around the 2,500-year return period, then gradually decreases to 10 for the rarest events. This non-monotonic trend is most likely due to a sampling error and is discussed in more detail in SM Sec. S5.2. Moreover, we hypothesize that the difference in variance speed-up for the two regions is due to regional differences in AI emulator weather forecast skill, potentially due to differences in soil moisture variability, as pointed out in SM Fig. S8.
@@
Figure 4: Composite maps of heatwaves over France with return periods exceeding 100 years. Events are defined by Eq. (1) with $L=$ 3 days. From top to bottom, the rows show daily mean Z500 anomaly composites three days before the heatwave onset, the $L$-day average Z500 anomaly composites during the heatwave, and the $L$-day average T2m anomaly composites during the heatwave.
A major advantage of AI+RES over the EVT baseline is that it generates full physical trajectories leading to the rare event of interest. This is particularly valuable for investigating the dynamics of the rare event and its precursors [18].
In Fig. 4, we illustrate this capability by presenting composite maps associated with heatwaves over France with a return period longer than 100 years from the ground-truth DNS (N=50, [^0] N=50{,}000), samples generated with AI+RES ($N=400$), and samples generated from the DNS with the same ensemble ($N=400$). A strong precursor pattern emerges in the ground-truth DNS three days before the heatwave onset, marked by a pronounced positive 500-hPa geopotential height anomaly over the British Isles, flanked by a negative anomaly over the Labrador Sea. A weaker negative anomaly off the coast of Portugal suggests potential influence from a cut-off low, while a weak negative anomaly is also visible over Eastern Europe. At the heatwave peak, this synoptic pattern shifts: the high-pressure anomaly migrates eastward, the Labrador Sea anomaly propagates into the North Atlantic, and the Eastern European anomaly intensifies. The AI+RES composite shows similar patterns and captures the correct evolution, although it is slightly noisier than the ground truth due to fewer samples. In contrast, the composite from the DNS with the same computational budget as AI+RES is significantly noisier and contains incorrect physical patterns. For example, the 3-day-lagged Z500 anomaly north of Scandinavia in this composite is a spurious teleconnection likely due to sampling error.
Similarly, SM Fig. S13 shows that AI+RES generates physically realistic samples over Chicago that match those from ground truth, while DNS with the same computational budget produces samples with spurious features, obscuring the underlying dynamics.